A square plate of copper with 55.0 cm sides has no net charge and is placed in a region of uniform electric field of 83.0 kN/C directed perpendicularly to the plate.

(a) Find the charge density of each face of the plate.

(b) Find the total charge on each face.
Answer #1

a) E_ext = 2E = (local surface charge density)/eps_0

so, (local surface charge density) = E_ext*eps_0.
83.0 kN/C = 83x10^3
(83x10^3) * (8.85x10^-12) = 7.34x10^-7 or 734 nC/m^2

the answer will neg and pos... so you will have two answers... -734, 734

b) for each plate ----> multiply the answers by the sides squared.
-734 * 0.55^2 = - 222.035
734 *0 .55^2 = 222.035